The titles LaTeX package

title macros (Frankenstein’s references)

              Matt Swift <swift@alum.mit.edu>

Version: 1.2 Date: 2001/08/31

Documentation revision: 2001/08/31

Abstract

The titles package defines macros that typeset the titles of books, journals, etc. and handle following spacing and punctuation intelligently, based on context. These are useful for bibliographic databases, for example. Also defined is other markup like \word, \defn, \phrase, etc.

Contents

I Discussion 3

1 Options 3

2 Words and phrases 3

3 Titles 4

4 Programmer’s interface 6

4.1 Limitations of Wrapquotes and friends . . . . . . . . . . . . . . . . 7 4.1.1 Nesting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.1.2 Italic corrections . . . . . . . . . . . . . . . . . . . . . . . . 7 4.1.3 A slight bug . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

II Implementation 9

5 Version control 9

6 Requirements 9

7 Options 9

8 Wrapquotes 10 8.1 Titles that are Questions or Exclamations . . . . . . . . . . . . . . 10 8.2 Highlevel macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.3 Opening quotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 8.4 Closing macros that don’t suck . . . . . . . . . . . . . . . . . . . . 14 8.5 Closing macros that suck . . . . . . . . . . . . . . . . . . . . . . . 15 8.6 Looking ahead . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

9 Words and phrases 21

10 Titles 22

III Configuration 23

11 User Customization 23

IV Testing 24

11.1 Question and exclamation marks . . . . . . . . . . . . . . . . . . . 24 11.2 Plain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.3 Nested beginnings . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.4 Nested endings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.5 double and single nosuck . . . . . . . . . . . . . . . . . . . . . . . . 26

Part I

Discussion

1 Options

There are two package options, british and american, the default is american. They select the conventional way to use quotation marks: British style is use single quotes, and do not suck following period or comma inside; American style is to use double quotes and to suck following period or comma inside.

2 Words and phrases

\word Typeset a word or phrase referred to as a noun with \word {⟨word⟩}. The argument is not expected to contain punctuation.

     \word{Elephant} is such a silly word.

LOOKS LIKE:

Elephant is such a silly word.

\phrase Typeset a phrase used as a noun rather than direct quotation with \phrase {⟨phrase⟩}. The argument might well have punctuation, including final punctuation, which should not be considered to be punctuation of the containing sentence.

     The sentence \phrase{And stop calling me Shirley!} occurs
     twenty-seven times.

LOOKS LIKE:

The sentence ‘And stop calling me Shirley!’ occurs twenty-seven times.

\foreign Typeset a foreign word or phrase with \foreign {⟨foreign text⟩}.

     I couldn't think of the \foreign{mot juste} at the time.

LOOKS LIKE:

I couldn’t think of the mot juste at the time.

\foreignword Typeset a foreign word or phrase referred to as a noun with \foreignword {⟨foreign word⟩}.

     Only later did I realize that the right word was
     \foreignword{bouffon}.

LOOKS LIKE:

Only later did I realize that the right word was ‘bouffon’.

Warning: Notice that writing \foreign{\word{text}} or vice versa is not necessarily going to do the right thing. Suppose \foreign and \word were both set to \textitswitch (which are in fact the default settings below). Then \foreign{\word{text}} is going to cancel out and look just like the surrounding text. This is not the most intuitive fact, but it’s not worth it to try to make \foreign and \word smart enough to see each other inside themselves.

\term \term {⟨technical term⟩} typesets a techincal term in a different font. You might want to use this where a techincal term is first used, or defined. One could enhance this macro and \defn to help build an automatic glossary

 This sort of thing is called a \term{blibnil}.

LOOKS LIKE:

This sort of thing is called a blibnil.

\defn \defn {⟨definition⟩} typesets a definition, perhaps of a technical term. One could enhance this macro and \defn to help build an automatic glossary

We may describe a \term{blibnil} as \defn{a slibnil with
three arms}.

LOOKS LIKE:

We may describe a blibnil as a slibnil with three arms.

3 Titles

\book \book {⟨book title⟩} typesets a book title.

Some people find \book{Moby-Dick} dull, but I thought
it was exciting.

LOOKS LIKE:

Some people find Moby-Dick dull, but I thought it was exciting.

\journal \journal {⟨journal title⟩} typesets a journal title.

I liked it so much I started a scholarly journal called
\journal{The Melville Times} with the inheritance from
my grandmother.

LOOKS LIKE:

I liked it so much I started a scholarly journal called The Melville Times with the inheritance from my grandmother.

\music \music {⟨music title⟩} typesets a music title.

My journal didn't do very well; I moped around my office
and listened to Schubert's \music{Winterreise}.

LOOKS LIKE:

My journal didn’t do very well; I moped around my office and listened to Schubert’s Winterreise.

\article \article {⟨article title⟩} typesets a article title.

Then one day I received an article, \article{Pip and
the Milk of Human Kindness}, by express mail from Wales.

LOOKS LIKE:

Then one day I received an article, “Pip and the Milk of Human Kindness,” by express mail from Wales.

\poemtitle \poemtitle {⟨poem title⟩} typesets a poem title.

I then wrote my famous poem \poemtitle{Jump for Joy like
the Butterflies of Troy} in five minutes.

LOOKS LIKE:

I then wrote my famous poem “Jump for Joy like the Butterflies of Troy” in five minutes.

Sometimes longer poems are distinguished from shorter ones in type when they have been published separately as a book [FIX give reference]. This package defines a macro \longpoem in the configuration file in the following way:

\newlet\longpoem\textitswitch

\play \play {⟨play title⟩} typesets a play title.

To celebrate the popularity of the article, I took the
author to the theater to see the acclaimed play
\play{Grave in Waterloo}, starring Vincent Price.

LOOKS LIKE:

To celebrate the popularity of the article, I took the author to the theater to see the acclaimed play Grave in Waterloo, starring Vincent Price.

\craft \craft {⟨craft title⟩} typesets a title of a craft or ship.

With tears in my heart, I put the author on the \craft{HMS Shangrila}
bound for Wales.

LOOKS LIKE:

With tears in my heart, I put the author on the HMS Shangrila bound for Wales.

\species \species{⟨[ genus ] species [ subspecies ]⟩} typesets the Latin generic and/or specific names for an organism.

Lesson 1 Chicago Manual of Style specifies italic type. Genus names should be capitalized, and may be abbreviated on subsequent appearances with the initial letter. Following designations should be in roman. E.g., “var.” for “variant” following species name and “sp.” for “species” following genus name, meaning “any species in the genus.” §7.102–4

Higher groupings should be in capitalized roman. English derivatives of scientific names, e.g., amoeba, are lowercased. §7.105–6

To do: abbrevs category for genus/species and/or datemark for suffixes

Warning: Right now there is a small discrepancy between the behavior of \textitswitch and \Wrapquotes regarding what happens when followed by a command sequence such as \footnote. I hope to make these things completely parallel one day, but for now, realize that after using a titling macro that uses \Wrapquotes, you must use {} before any following command sequence that you want to immediately follow the title with no intervening space. The only case I can think of is \footnote. If you forget the {}, you will have an extra space after the title and before the footnotemark. The following example illustrates this behavior and contrasts it with \textitswitch:

\newabbrev\foo{Foo}
\book{Foo}\foo
\book{Foo} \foo
\book{Foo}\footnote{footie}
\book{Foo}{}\footnote{footie}
\poemtitle{Foo}\foo
\poemtitle{Foo} \foo
\poemtitle{Foo}\footnote{footie}
\poemtitle{Foo}{}\footnote{footie}
\poemtitle{Foo}.\footnote{footie}

LOOKS LIKE:

FooFoo Foo Foo Foo^{a} Foo^{b} “Foo” Foo “Foo” Foo “Foo” ^{c} “Foo”^{d} “Foo.”^{e}

afootie bfootie cfootie dfootie efootie

4 Programmer’s interface

\Wrapquotes \Wrapquotes {⟨text⟩} wraps ⟨text⟩ in quotes. Single quotes are used initially if

\WrapquotesNS
\WrapquotesIS
\WrapquotesNN
\WrapquotesIN
\WrapquotesSN
\WrapquotesDN
\WrapquotesSK

6 the singlequotes option is given to the package, and double quotes if no option or the doublequotes option is given to the package.

When quotation marks inserted by \Wrapquotes and friends are doubled up (this occurs sometimes when nesting them), a thinspace (\,) is inserted between the abutted quotes.

\Wrapquotes will be \let to one of the six macros \Wrapquotes ⟨XY ⟩.

In the two-letter suffix ⟨XY ⟩, first letter N means “normal” and I means “inverse.” These are macros that switch between single and double quotes when they nest: an inverse wrapquotes wraps with single quotes when a normal wrapquotes would wrap with double quotes, and vice versa. First letter S for “single” and D for “double” are for macros that always wrap with single or double quotes. Spacing and punctuation following the closing quotes are handled intelligently by macros with second letter S, which means means suck a following period or comma into the closing quote, that is, if what follows is a comma or period, it is pulled inside the quotes (following American practice). Second letter N means “nosuck,” that is, don’t suck. Second letter K means “kill”: the same as N but suppress the effect of any punctuation in the quoted argument on spacing that follows the closing quotes (i.e., execute \@, which sets the spacefactor to 1000). This is only useful in certain technical writing where punctuation in the quoted argument should not be considered puncutation of the containing sentence.

A space is inserted after the closing quotes unless what follows is in the set ;?:!-)]'\textquoteright{, in which case no space is inserted.FIX: that would

be \nospacelist
\IfQuestionOrExclamation      \IfQuestionOrExclamation {⟨text⟩}{⟨true⟩}{⟨false⟩} executes ⟨true⟩ clause

iff ⟨text⟩ ends with a question mark or an exclamation point; executes ⟨false⟩ clause otherwise.

4.1 Limitations of Wrapquotes and friends

4.1.1 Nesting

Warning: For proper nesting of \Wrapquotes and friends, user commands must be \let to \Wrapquotes or one of the six \Wrapquotes ⟨XY ⟩ commands instead of using a \def-like defining command. It’s OK to \let a user macro to something like \Wrapquotes which itself has been \let to one of the six \Wrapquotes ⟨XY ⟩ macros.

The user command which is \let to one of the \Wrapquotes commands must furthermore appear in the source. That is, it must not appear as the result of an expansion. Among other things, this means that nesting won’t work properly if you put \Wrapquotes into an abbrev (see the abbrevs package in the Frankenstein bundle).

For applications where nesting will not occur, there should be nothing to worry about.

4.1.2 Italic corrections

Warning: The question of when to insert an italic correction is not nearly as simple as it might seem. I cannot figure good rules which cover all cases, and I do not trust the behavior of the kernel’s macros as a guide. So I can not tell you whether this package handles italic corrections properly. If you discover behavior which you think is wrong, please let me know with an example and an argument.

4.1.3 A slight bug

Warning: Right now there is a small bug in cases where closing quotes fall at the end of italic text, such as

          \normalfont
          \book{My love of \poemtitle{Daffodils}}, by H.~Moneysworth.

LOOKS LIKE:

My love of “Daffodils,” by H. Moneysworth.

These cases loose because the closing quotation marks and any sucked-in punctuation are going to be in roman, not italic, or italic, not roman. Only the more obsessive will notice this flaw. I’m sure I will come up with a way to handle this for a future version of this package.

Part II

Implementation

5 Version control

\fileinfo \DoXUsepackagE \HaveECitationS \fileversion \filedate \docdate \PPOptArg These definitions must be the first ones in the file.
 1 \def\fileinfo{title macros (Frankenstein's references)}
 2 \def\DoXPackageS {}
 3 \def\initelyHavECitationS {}
 4 \def\fileversion{v1.2}
 5 \def\filedate{2001/08/31}
 6 \def\docdate{2001/08/31}
 7 \edef\PPOptArg {%
 8 \filedate\space \fileversion\space \fileinfo
 9 }

If we’re loading this file from a \ProcessDTXFile command (see the compsci package), then \JusTLoaDInformatioN will be defined; othewise we assume it is not (that’s why the FunkY NamE).

If we’re loading from \ProcessDTXFile, we want to load the packages listed in \DoXPackageS (needed to typeset the documentation for this file) and then bail out. Otherwise, we’re using this file in a normal way as a package, so do nothing. \DoXPackageS, if there are any, are declared in the dtx file, and, if you’re reading the typeset documentation of this package, would appear just above. (It’s OK to call \usepackage with an empty argument or \relax, by the way.)

10 \makeatletter% A special comment to help create bst files.  Don't change!
11 \@ifundefined{JusTLoaDInformatioN} {%
12 }{% ELSE (we know the compsci package is already loaded, too)
13 \UndefineCS\JusTLoaDInformatioN
14 \SaveDoXVarS
15 \eExpand\csname DoXPackageS\endcsname\In {%use \csname in case it's undefined
16  \usepackage{#1}%
17 }%
18 \RestoreDoXVarS
19 \makeatother
20 \endinput
21 }% A special comment to help create bst files.  Don't change!

Now we check for L^{A}TEX2e and declare the LaTeX package.

22 \NeedsTeXFormat{LaTeX2e}
23 \ProvidesPackage{titles}[\PPOptArg]

6 Requirements

24 \RequirePackage{moredefs,slemph}

7 Options

\ti@domelater
25 \ReserveCS\ti@domelater

26 \DeclareOption{british} {%
27 \def\ti@domelater {%
28  \let\Wrapquotes\WrapquotesNN
29  \@doublequotes@false

30 } 31 }

32 \DeclareOption{american} {%
33 \def\ti@domelater {%
34  \let\Wrapquotes\WrapquotesNS
35  \@doublequotes@true

36 } 37 }

38 \ExecuteOptions{american}
39 \ProcessOptions

8 Wrapquotes

Here we go! This is not a picnic, so leave your jelly jar home.

8.1 Titles that are Questions or Exclamations

\IfQuestionOrExclamation
\ti@checkfor@q
\ti@checkfor@e
\ti@prev
\ti@prev@prev
\@ti@sw@true
\@ti@sw@false
\if@ti@sw@

40 \newcommand\IfQuestionOrExclamation [1] {%
41 \@tempswafalse
42 \ti@checkfor@q #1?\@nil
43 \ti@checkfor@e #1!\@nil
44 \if@tempswa
45  \expandafter\@firstoftwo
46 \else
47  \expandafter\@secondoftwo
48 \fi

49 }

The large majority of titles will not contain a question mark or exclamation point. The large majority of those that do will have a single mark or point at the end. We could (I think) use a simpler check that processed all titles by looping through to examine the end, but a slightly more complicated check will handle the majority of cases very quickly (and at a constant speed, rather than proportional to title length) and not greatly slow down processing the remaining two unusual cases. We divide our argument (with an extra question mark tacked onto the end) into what’s before the first question mark and what’s after it. Then we examine what’s after it and interpret the results thus:

empty no question mark in title

question mark title ends with question mark (and there are no other question marks)

text ending with one question mark a question mark occurs in the title, but not at the end

text ending with two question marks title ends with a question mark (and there is a previous question mark)

We set switch a to true if the title ends with a question mark.

50 \newboolean{@ti@sw@}
51 \ReserveCS\ti@prev
52 \ReserveCS\ti@prev@prev
53 \NewName{ti@checkfor@q} {#1?#2\@nil} {%
54 \def\sc@t@a{#2}%
55 \def\sc@t@b{?}%
56 \ifx\sc@t@a\ShortEmpty
57 \else
58  \ifx\sc@t@a\sc@t@b
59  \@tempswatrue
60  \else

We use a loop to whittle down #2 until \ti@prev contains the last character and \ti@prev@prev contains the second-to-last. We know that \ti@prev is going to be a question mark. Iff \ti@prev@prev is a question mark, we are in the final case above.

61  \let\ti@prev\sc@t@a
62  \let\ti@prev@prev\sc@t@a
63  \@ti@sw@true
64  \@whilesw \if@ti@sw@ \fi {%
65  \ifx\sc@t@a\ShortEmpty
66  \@ti@sw@false
67  \else
68  \let\ti@prev@prev\ti@prev
69  \let\ti@prev\sc@t@a
70  \edef\sc@t@a{\E@cdr\sc@t@a\@nil}%
71  \fi
72  }%
73  \edef\ti@prev@prev{\E@car\ti@prev@prev\@nil}%
74  \ifx\ti@prev@prev\sc@t@b
75  \@tempswatrue
76  \fi
77  \fi
78 \fi

79 }

Exact same logic applies to exclamation points.

80 \NewName{ti@checkfor@e} {#1!#2\@nil} {%
81 \def\sc@t@a{#2}%
82 \def\sc@t@b{!}%
83 \ifx\sc@t@a\ShortEmpty
84 \else
85  \ifx\sc@t@a\sc@t@b
86  \@tempswatrue
87  \else
88  \let\ti@prev\sc@t@a
89  \let\ti@prev@prev\sc@t@a
90  \@ti@sw@true
91  \@whilesw \if@ti@sw@ \fi {%
92  \ifx\sc@t@a\ShortEmpty
93  \@ti@sw@false
94  \else
95  \let\ti@prev@prev\ti@prev

 96  \let\ti@prev\sc@t@a
 97  \edef\sc@t@a{\E@cdr\sc@t@a\@nil}%
 98  \fi
 99  }%
100  \edef\ti@prev@prev{\E@car\ti@prev@prev\@nil}%
101  \ifx\ti@prev@prev\sc@t@b
102  \@tempswatrue
103  \fi
104  \fi
105 \fi

106 }

8.2 Highlevel macros

\ti@wrapquotes@suck \ti@wrapquotes@nosuck These two are the top-level internal macros, and they are pretty sane. One sucks in a following period or comma, the other does not. \ti@wrapquotes@suck does not suck, however, when the title ends in a question or exclamation point.

The group here is necessary to scope the \@doublequotes@ boolean.

107 \newcommand*\ti@wrapquotes@suck [1] {%
108 %  \DTypeout{top of wrapquotes@suck}%
109 \IfQuestionOrExclamation {#1} {%
110  \ti@wrapquotes@nosuck{#1}%
111  }{% ELSE
112 %    \DTypeout{top of wrapquotes@suck ELSE}%
113  \begingroup
114  \if@doublequotes@
115 %        \DTypeout{double true in suck}%
116  \@doublequotes@false
117  \def\sc@t@a {\ti@open@double #1\ti@close@double@suck}%
118  \else
119 %        \DTypeout{double false in suck}%
120  \@doublequotes@true
121  \def\sc@t@a {\ti@open@single #1\ti@close@single@suck}%
122  \fi
123  \sc@t@a
124  \endgroup

125 }% 126 }

127 \newcommand*\ti@wrapquotes@nosuck [1] {%
128 \begingroup
129  \if@doublequotes@
130 %      \DTypeout{double true in nosuck}%
131  \@doublequotes@false
132  \def\sc@t@a {\ti@open@double #1\ti@close@double@nosuck}%
133  \else
134 %      \DTypeout{double false in nosuck}%
135  \@doublequotes@true
136  \def\sc@t@a {\ti@open@single #1\ti@close@single@nosuck}%
137  \fi
138  \sc@t@a
139 \endgroup

140 }

\WrapquotesNS \WrapquotesIS \WrapquotesNN \WrapquotesIN \WrapquotesSN \WrapquotesDN \WrapquotesSK \Wrapquotes \if@doublequotes@ \@doublequotes@true \@doublequotes@false Now we can define the secondary programmers’ macros.

We simply reserve \Wrapquotes here, and assign it in the user options section above.

141 \newboolean{@doublequotes@}
142 \newcommand*\WrapquotesNS {%
143 %  \DTypeout{starting wrapquotes NS}%
144 \ti@wrapquotes@suck

145 }

146 \newcommand*\WrapquotesIS {%
147 %  \DTypeout{starting wrapquotes IS}%
148 \ToggleBoolean{@doublequotes@}%
149 \ti@wrapquotes@suck

150 }

151 \newcommand*\WrapquotesNN {%
152 %  \DTypeout{starting wrapquotes NN}%
153 \ti@wrapquotes@nosuck

154 }

155 \newcommand*\WrapquotesIN {%
156 %  \DTypeout{starting wrapquotes IN}%
157 \ToggleBoolean{@doublequotes@}%
158 \ti@wrapquotes@nosuck

159 }

160 \newcommand*\WrapquotesSN [1] {%
161 %  \DTypeout{starting wrapquotes SN}%
162 \begingroup
163  \ti@open@single #1\ti@close@single@nosuck
164 \endgroup

165 }

166 \newcommand*\WrapquotesDN [1] {%
167 %  \DTypeout{starting wrapquotes DN}%
168 \begingroup
169  \ti@open@double #1\ti@close@double@nosuck
170 \endgroup

171 }

172 \newcommand*\WrapquotesSK [1] {% FIX: test
173 %  \DTypeout{starting wrapquotes SK}%
174 \begingroup
175  \ti@open@single #1\ti@close@single@nosuck\@%
176 \endgroup

177 }

178 \ReserveCS\Wrapquotes
179 \ti@domelater

8.3 Opening quotes

\ti@open@double \ti@open@single \ti@openquote We start by putting an opening mark in scratch f with a global definition.

I can’t remember why it’s global. In the macros that close quotes, we want to keep that information around past a group end because we’re using \aftergroup, but that doesn’t seem to apply for opening them. Best not to change what’s not broke, however.

180 \newcommand\ti@open@double {%
181 \gdef\sc@t@f {\textquotedblleft}%
182 \ti@openquote

183 }

184 \newcommand\ti@open@single {%
185 \gdef\sc@t@f {\textquoteleft}%
186 \ti@openquote

187 }

Then we look ahead with scratch a. We are looking ahead at the first character of the contents of the \Wrapquotes.

188 \newcommand\ti@openquote {%
189 \futurelet\sc@t@a\ti@@openquote

190 }

Insert the opening mark. Then, if we are about to open another quote, insert the space appropriate to separate contiguous quotation marks.

191 \newcommand\ti@@openquote {%
192 \sc@t@f
193 \ifx\sc@t@a\WrapquotesNS
194 %    \DTypeout{Quotation marks are doubled up (next is NS); inserting padding.}%
195  \,%
196 \else \ifx\sc@t@a\WrapquotesNN
197 %    \DTypeout{Quotation marks are doubled up (next is NN); inserting padding.}%
198  \,%
199 \else \ifx\sc@t@a\WrapquotesIN
200 %    \DTypeout{Quotation marks are doubled up (next is IN); inserting padding.}%
201  \,%
202 \else \ifx\sc@t@a\WrapquotesIS
203 %    \DTypeout{Quotation marks are doubled up (next is IS); inserting padding.}%
204  \,%
205 \else \ifx\sc@t@a\WrapquotesSN
206 %    \DTypeout{Quotation marks are doubled up (next is SN); inserting padding.}%
207  \,%
208 \else \ifx\sc@t@a\WrapquotesDN
209 %    \DTypeout{Quotation marks are doubled up (next is DN); inserting padding.}%
210  \,%
211 \else \ifx\sc@t@a\WrapquotesSK
212 %    \DTypeout{Quotation marks are doubled up (next is SK); inserting padding.}%
213  \,%
214 \else
215 \fi \fi \fi \fi \fi \fi \fi

216 }

8.4 Closing macros that don’t suck

This case that doesn’t suck is easier, so we warm up with it.

\ti@close@single@nosuck
\ti@close@double@nosuck
\ti@close@single@@nosuck
\ti@close@double@@nosuck

217 \newcommand*\ti@close@single@nosuck {%
218 \aftergroup\ti@close@single@@nosuck

219 }

220 \newcommand*\ti@close@double@nosuck {%
221 \aftergroup\ti@close@double@@nosuck

222 }

223 \newcommand*\ti@close@single@@nosuck {%
224 \gdef\sc@t@f {\textquoteright}%
225 \ti@close@quote@nosuck

227 \newcommand*\ti@close@double@@nosuck {%
228 \gdef\sc@t@f {\textquotedblright}%
229 \ti@close@quote@nosuck

230 }

\ti@close@quote@nosuck \if@look@nosuck@ \@look@nosuck@true \@look@nosuck@false To do: Document this flag. It’s a hack, we must set it before each call to \ti@

q@ifnextcharin I think. What it stands for is something like the presence of the

tokens \ti@close@single@nosuck and \ti@close@double@nosuck in the list

of chars to look for, but since they aren’t really chars they can’t go in the list, so instead we set the flag. Somewhat cleaner would be putting a flag char in the list, but I can’t think of what char I could safely use.

231 \newboolean{@look@nosuck@}
232 \@look@nosuck@false
233 \newcommand\ti@close@quote@nosuck {%
234 %  \DTypeout{Starting ti@close@quote@nosuck}%
235 \@look@nosuck@true

FIX Aha, but here is a good reason to leave in ., in our substitute for

\nospacelist.
236 \expandafter \ti@q@ifnextcharin \expandafter {\nospacelist} {%
237 %      \DTypeout{Found a nosuck no-spacer.  C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
238  \sc@t@f
239  }{% ELSE
240 %      \DTypeout{Found a nosuck spacer. C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
241  \sc@t@f\space

242 }% 243 }

8.5 Closing macros that suck

\ti@close@double@suck \ti@close@single@suck We need to look ahead beyond the \endgroup that ends \Wrapquotes??. The

lookahead mechanism that gets invoked in scratch a below could handle looking past the \endgroup, but I think it is more efficient to skip it by using \aftergroup.

244 \newcommand\ti@close@double@suck {%
245 \aftergroup\ti@close@double@@suck

246 }

247 \newcommand\ti@close@single@suck {%
248 \aftergroup\ti@close@single@@suck

249 }

\ti@close@double@@suck \ti@close@single@@suck This part isn’t so bad yet. To close the quotes, we again start with the closing mark in scratch f, with a global definition.

250 \newcommand\ti@close@double@@suck {%
251 \gdef\sc@t@f {\textquotedblright}%
252 \ti@close@quote@suck

253 }

254 \newcommand\ti@close@single@@suck {%
255 \gdef\sc@t@f {\textquoteright}%
256 \ti@close@quote@suck

257 }

\nospacelist Put these in the order of their frequency. Anything in \nocorrlist should also be in here, most likely. I’m putting in \@xobeysp because it’s in the xspace package, but I can’t tell you when it would come up.

258 \requirecommand\nospacelist {%
259 ,.':;?-/\slash~!)]\bgroup\egroup\@sptoken\ \space\/\@xobeysp

260 }

\ti@close@quote@suck Then we use \ti@q@ifnextcharin to look as far ahead as necessary for a significant character. The latest significant character found is available in scratch c. The work of handling all the cases of what we might find while looking ahead

is divided up between \ti@close@quote@suck and \ti@q@ifnextcharin. \ti@
close@quote@suck handles the last step in the process, and \ti@q@ifnextcharin

handles all the steps up to the last.

Here is what \ti@close@quote@suck does, in English. If we find a comma or period, we put it inside the closing quote, and gobble the one we found. That is, we print out scratch c, then scratch f, then gobble a character. If we find something in the set given in \nospacelist, do not leave a space after the closing mark. That is, just print out scratch f. If we find something else, we leave a space after the closing mark. That is, print scratch f and a space.

261 \newcommand\ti@close@quote@suck {%
262 %  \DTypeout{Starting ti@close@quote@suck}%
263 \@look@nosuck@false
264 \ti@q@ifnextcharin {.,} {%
265 %      \DTypeout{Found a comma or period. C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
266  \sc@t@c\sc@t@f\DGobbleM % This gobbles the original punctuation.
267  }{% ELSE
268 %    \DTypeout {Before second ti@qifnextcharin.  C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
269  \@look@nosuck@true

To do: Using \nospacelist is inefficient here, since some of the cases, namely ,.\@sptoken}, are never going to be there and shouldn’t be checked for, since they are passed over by \ti@q@ifnextcharin before the list is compared. But it would be good to have this parallelism between abbrevs and titles.

270  \expandafter \ti@q@ifnextcharin \expandafter {\nospacelist} {%
271 %        \DTypeout{Found a suck no-spacer.  C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
272  \sc@t@f
273  }{% ELSE
274 %      \DTypeout{Found a suck spacer. C=[\meaning\sc@t@c] F=[\meaning\sc@t@f]}%
275  \sc@t@f\space

276 }% 277 }% 278 }

8.6 Looking ahead

Now things are getting fun.

\ti@q@ifnextcharin \ti@q@check \ti@q@ifnch \ti@q@@ifnch These macros are modeled after the definition of \@ifnextchar which skips spaces.

While looking ahead for the next significant character, these macros skip spaces,

\egroup, \endgroup, \check@icr, \ti@close@double and \ti@close@single

while doing the right thing after each.

The first argument should be a list of tokens. If the next significant char is in the list, then the true clause is executed, otherwise the false clause is executed. The next significant char is left in scratch c so it can be accessed by the clauses.

The three arguments to \ti@q@ifnextcharin are saved in global variables because while looking ahead we must continue past the ends of groups.

FIX Not sure I need gdef for scratch e.

279 \newcommand\ti@q@ifnextcharin [3] {% args: charlist true false
280 %  \sc@toks@a{#1}%
281 %  \DTypeout{charlist unexpanded is =[\the\sc@toks@a]}%
282 \gdef\sc@t@e {#1}%
283 \gdef\sc@t@a {#2}%
284 \gdef\sc@t@b {#3}%
285 \ti@q@check

286 }

Having saved the arguments, we look ahead with scratch c. This step is not in the macro above so that we can jump back to \ti@q@check whenever we want to look ahead another character.

287 \newcommand\ti@q@check {%
288 \futurelet\sc@t@c\ti@q@ifnch

289 }

Scratch c contains the current char. Scratch d is the action to take at the end of this macro. We attempt to order these possibilities to make \Wrapquotes most efficient, though it is a guess which items will be encountered most frequently.

The actions taken for each of the possibilities are the following:

\ifvmode Assume that the \Wrapquotes was the argument of a \TextFontCommand

from certain L^{A}TEX kernels. Gobble three more tokens expected to follow the \ifvmode, execute them, and continue on to look ahead another character. See documentation of \ti@q@handle@ifvmode for more details.

\check@icr This means the \Wrapquotes was the argument of a \TextFontCommand.

Gobble the \check@icr and look ahead another character after we exit the group that the \TextFontCommand has given us.

\endgroup and } Pass right by an \endgroup or } and look ahead another char.

\@sptoken (a non-explicit space) Handle a non-explicit space by calling \ti@ q@handle@space, which gobbles the space and looks ahead another char. When the user or a macro has followed the titles with an explicit space such as a tie, or the \␣ or \space macros, we do nothing and let this be caught by the comparison to the tokens in the argument of \ti@q@ifnextcharin.

\ti@close@double@suck, \ti@close@single@suck
\ti@close@double@nosuck, and \ti@close@single@nosuck We are in a nested
     \Wrapquotes. Call \ti@q@handle@single/double@suck/nosuck as appro-

priate, which gobbles the closequotes token, adds properly-padded closing quotes to scratch f, and then goes on to look ahead another character.

The lookahead process stops when it finds something not on this list. Then it compares what it found to the list of characters given to \ti@q@ifnextcharin and executes the true or false clause as appropriate.

First we have to handle the case of finding an \ifvmode. We can’t bundle this test in with the tests for other tokens, so it gets its own macro, \ti@q@handle@ ifvmode, which see for details. The remaining cases are handled in \ti@q@@ifnch.

290 \newcommand\ti@q@ifnch {%
291 %  \DTypeout{The lookahead in ti@q@ifnch: [\meaning\sc@t@c]}%
292 \ifx\sc@t@c\ifvmode
293  \let\sc@t@d\ti@q@handle@ifvmode
294 \else
295  \let\sc@t@d\ti@q@@ifnch
296 \fi
297 \sc@t@d

298 }

299 \newcommand\ti@q@@ifnch {%
300 %  \DTypeout{entering ti@q@@ifnch}%
301 %  \expandafter\sc@toks@a\expandafter{\sc@t@c}%
302 %  \DTypeout{ti@q@@ifnch: C expanded once is =[\the\sc@toks@a]}%
303 \ifx\sc@t@c\check@icr
304 %    \DTypeout{Handling check@icr}%
305  \defcommand\sc@t@d [1] {%
306 %      \DTypeout{check@icr handler: gobbling [\meaning ##1]}%
307  ##1\aftergroup\ti@q@check

308 }%

309 \else \ifx\sc@t@c\endgroup
310 %    \DTypeout{Handling endgroup}%
311  \def\sc@t@d {\aftergroup\ti@q@check}%
312 \else \ifx\sc@t@c\@sptoken
313 %    \DTypeout{Handling space}%
314  \let\sc@t@d\ti@q@handle@space
315 \else \ifx\sc@t@c\egroup
316 %    \DTypeout{Handling egroup}%
317  \def\sc@t@d {\aftergroup\ti@q@check}%
318 \else \ifx\sc@t@c\ti@close@double@suck
319 %    \DTypeout{Handling ti@close@double@suck}%
320  \let\sc@t@d\ti@q@handle@double@suck
321 \else \ifx\sc@t@c\ti@close@single@suck
322 %    \DTypeout{Handling ti@close@single@suck}%
323  \let\sc@t@d\ti@q@handle@single@suck
324 \else \ifx\sc@t@c\ti@close@double@nosuck
325 %    \DTypeout{Handling ti@close@double@nosuck}%
326  \let\sc@t@d\ti@q@handle@double@nosuck
327 \else \ifx\sc@t@c\ti@close@single@nosuck
328 %    \DTypeout{Handling ti@close@single@nosuck}%
329  \let\sc@t@d\ti@q@handle@single@nosuck
330 \else

We’ve handled all the lookahead cases, so now we are left with the simple comparison of the next char with the charlist.

331  \@tempswafalse
332  \expandafter   \@tfor
333  \expandafter \sc@t@g
334  \expandafter :%
335  \expandafter =%
336  \sc@t@e
337  \do {%

338  \expandafter\ifx\sc@t@g\sc@t@c
339 %                  \DTypeout{We have a match of [\meaning\sc@t@c]
340 %                           with [\expandafter\meaning\sc@t@g]}%
341  \@tempswatrue
342  \@break@tfor
343  \else
344 %                  \DTypeout{We have NO match between [\meaning\sc@t@c]
345 %                           with [\meaning\sc@t@g]}%
346  \fi

347 }%

348  \if@tempswa
349 %                \DTypeout{Choosing true clause [\meaning\sc@t@a]}%
350  \let\sc@t@d\sc@t@a % the ``true'' clause
351  \else
352 %                \DTypeout{Choosing false clause [\meaning\sc@t@b]}%
353  \let\sc@t@d\sc@t@b % the ``false'' clause
354  \fi
355 \fi \fi \fi \fi \fi \fi \fi \fi
356 %  \DTypeout{About to fall out of ti@q@@ifnch and do this [\meaning\sc@t@d]}%
357 \sc@t@d

358 }

\ti@q@handle@ifvmode This is in its own macro for clarity and to avoid problems with skipping over

clauses.

\ti@q@ifnch has to take two different kinds of L^{A}TEX kernel into account. The 1996/12/01 and 1997/06/01 kernels used a different definition of

\DeclareTextFontCommand:
359 %\def \DeclareTextFontCommand #1#2{%
360 %  \DeclareRobustCommand#1[1] {%
361 %    \ifmmode
362 %      \nfss@text{#2##1}%
363 %    \else
364 %      \leavevmode
365 %      {\text@command{##1}%
366 %       #2\check@icl ##1\ifvmode\else\check@icr\fi
367 %       \expandafter}%
368 %    \fi
369 %  }%

370 %}

All other kernels leave out the check for vertical mode (kernels from 1997/12/01 include it when necessary inside \check@icr). The macro \ti@q@ifnch, which will be called immediately before this point of difference, handles both cases by looking for both \ifvmode and \check@icr. For the history, see L^{A}TEX bug report 2646.

The check for \ifvmode must not be part of a nested conditional. TEX can’t match \ifs with \fis properly when you nest tests for \if-type tokens. See p. 211 of the TEXbook.

When we encounter an \ifvmode, we must assume we are inside a TextFont-Command declared by one of the two kernel versions mentioned above. If not, we are in an unknown situation and we will bomb. Since the error message in this case won’t be helpful, we warn the user in the log file. We use scratch d to gobble both the \ifvmode and what we expect will follow the \ifvmode, namely \else\check@icr\fi. After swallowing those, we reissue those same commands and then proceed with our lookahead. We want to issue those commands, which conditionally introduce an italic correction, before looking further ahead.

371 \newcommand\ti@q@handle@ifvmode {%
372 %  \DTypeout{Handling ifvmode}%
373 \FrankenInfo{titles}
374  {Handling an \string\ifvmode\space following a title.\MessageBreak
375  If you now get an error that \string\sc@t@d\space does not\MessageBreak
376  match its definition, this \string\ifvmode\space is\MessageBreak
377  unexpected}%
378 \DefName{sc@t@d} {\ifvmode\else\check@icr\fi} {%
379  \ifvmode
380  \else
381  \check@icr
382  \fi
383  \aftergroup\ti@q@check

384 }%

385 \sc@t@d

386 }

\ti@q@handle@space Handle the case of a following space: gobble the space and call \ti@q@check.

This little bit of trickery sneaks a space in as the \def template, thereby causing a space following \ti@q@handle@space to get gobbled. We use the control character \: and restore its value.

387 \ReserveCS\ti@q@handle@space
388 \let\sc@t@a\:
389 \def\:{\ti@q@handle@space} \expandafter\def\: {\ti@q@check}
390 \let\:\sc@t@a

\ti@q@handle@single@suck \ti@q@handle@double@suck Handle the single and double sucking cases: gobble the closequotes token with a \def template, add some stuff to scratch f and call \ti@q@check. These are put in their own macros only to avoid clutter above.

391 \newcommand*\ti@q@handle@double@suck [1] {%
392 %  \DTypeout{handle double suck: gobbling [\meaning#1]}%
393 %  \DTypeout{scratch f before: [\meaning\sc@t@f]}%
394 \g@addto@macro\sc@t@f {\,\textquotedblright}%
395 %  \DTypeout{scratch f after: [\meaning\sc@t@f]}%
396 \ti@q@check

397 }

398 \newcommand*\ti@q@handle@single@suck [1] {%
399 %  \DTypeout{handle single suck: gobbling [\meaning#1]}%
400 %  \DTypeout{scratch f before: [\meaning\sc@t@f]}%
401 \g@addto@macro\sc@t@f {\,\textquoteright}%
402 %  \DTypeout{scratch f after: [\meaning\sc@t@f]}%
403 \ti@q@check

404 }

\ti@q@handle@single@nosuck \ti@q@handle@double@nosuck Handle the single and double nosucking cases. Add inter-quote space to scratch f and exit \ti@q@@ifnch with true or false depending on whether we were looking for it. We had to do it this way instead of the normal \if test above at the end

 of \ti@q@@ifnch because \ti@close@double@nosuck is more than one character

long.

405 \newcommand*\ti@q@handle@double@nosuck [1] {%
406 % \DTypeout{handle double nosuck: gobbling [\meaning#1]}%
407 \if@look@nosuck@
408 %      \DTypeout{And we're looking for \string\ti@close@double@nosuck.}%
409  \g@addto@macro\sc@t@f {\,\textquotedblright}%
410 %      \DTypeout{After adding padding, F=[\meaning\sc@t@f]}%
411  \let\sc@t@d\sc@t@a % the ``true'' clause of ti@q@ifnextcharin
412 \else
413 %    \DTypeout{But we're not looking for \string\ti@close@double@nosuck.}%
414 %    \DTypeout{F is unchanged, F=[\meaning\sc@t@f]}%
415  \let\sc@t@d\sc@t@b % the ``false'' clause of ti@q@ifnextcharin
416 \fi
417 \ti@q@check

418 }

419 \newcommand*\ti@q@handle@single@nosuck [1] {%
420 % \DTypeout{handle single nosuck: gobbling [\meaning#1]}%
421 \if@look@nosuck@
422 %      \DTypeout{And we're looking for \string\ti@close@single@nosuck.}%
423  \g@addto@macro\sc@t@f {\,\textquoteright}%
424 %      \DTypeout{After adding padding, F=[\meaning\sc@t@f]}%
425  \let\sc@t@d\sc@t@a % the ``true'' clause of ti@q@ifnextcharin
426 \else
427 %    \DTypeout{But we're not looking for \string\ti@close@single@nosuck.}%
428 %    \DTypeout{F is unchanged, F=[\meaning\sc@t@f]}%
429  \let\sc@t@d\sc@t@b % the ``false'' clause of ti@q@ifnextcharin
430 \fi
431 \ti@q@check

432 }

9 Words and phrases

\word
\foreign
\foreignword
\phrase
\term
\defn

433 \newlet\word\textitswitch
434 \newlet\foreign\textitswitch

To do: \phrase is the result of expansion here: what effect will this have on its proper nesting, and is this something to worry about?

435 \newcommand\foreignword [1] {%
436 \phrase{\word{#1}}%

437 }

The \@ cancels the effect on spacing of any final punctuation in the argument.

438 % \newlet\phrase\WrapquotesSK -- whoops, doesn't work as intended,
439 % \phrase{foo}s puts a space before the following `s'

440

441 % old definition:
442 \newcommand\phrase [1] {%
443 \textquoteleft #1\textquoteright\@%

444 }

445 \newlet\term\textitswitch
446 \newlet\defn\textslswitch

10 Titles

\book
\journal
\music
\article
\storytitle
\poemtitle
\play
\craft
\species

447 \newlet\book\textitswitch
448 \newlet\journal\textitswitch
449 \newlet\music\textitswitch
450 \newlet\article\Wrapquotes
451 \newlet\storytitle\Wrapquotes
452 \newlet\poemtitle\Wrapquotes
453 \newlet\play\textitswitch % \manualref{7.145}
454 \newlet\craft\textitswitch
455 \newlet\species\textitswitch

Part III

Configuration

User alterations and additions and package testing are in a configuration file.

1 \InputIfFileExists{titles.cfg}{}{}

The contents of the distributed configuration file are below.

2 \def\fileinfo{titles package configuration}
3 \def\fileversion{v1.4}
4 \def\filedate{2001/08/31}
5 \def\docdate{2001/08/31}
6 \ProvidesFile{titles.cfg}

11 User Customization

Put your own alterations and additions here. For example.

7 % \let\word\textslswitch
8 \newlet\longpoem\textitswitch
9 \newlet\film\textitswitch
10 \newlet\essaytitle\Wrapquotes
11 \newlet\chaptertitle\Wrapquotes

Part IV

Testing

11.1 Question and exclamation marks

Test string: [ Title ] Result: Declarative Test string: [ Title? ] Result: Question or Exclamation Test string: [ Title! ] Result: Question or Exclamation Test string: [ Title?? ] Result: Question or Exclamation Test string: [ Title!! ] Result: Question or Exclamation Test string: [ Title? Title ] Result: Declarative Test string: [ Title! Title ] Result: Declarative Test string: [ Title!? ] Result: Question or Exclamation Test string: [ Title?! ] Result: Question or Exclamation Test string: [ Title? Title? ] Result: Question or Exclamation Test string: [ Title? Title! ] Result: Question or Exclamation Test string: [ Title! Title? ] Result: Question or Exclamation Test string: [ Title?? Title ] Result: Declarative Test string: [ Title!! Title ] Result: Declarative

11.2 Plain

Book Title. Test. Book Title, test. Book Title; test. Book Title test. Play Title. Test. “Play Title.” Test. Play Title, test. “Play Title,” test. Play Title; test. “Play Title”; test. Play Title test. “Play Title” test. title tie “title” tie title explicit space “title” explicit space

title \space

“title” \space title/slash “title”/slash title italcorr “title” italcorr title xobey “title” xobey

11.3 Nested beginnings

Book Title begins first book title and outside. Book Title, begins first book title, and outside. Book Title. begins first book title. and outside. Book Title; begins first book title; and outside. Play Title begins first book title and outside. Play Title, begins first book title, and outside. Play Title. begins first book title. and outside. Play Title; begins first book title; and outside. Book Title begins first play title and outside. Book Title, begins first play title, and outside. Book Title. begins first play title. and outside. Book Title; begins first play title; and outside. Play Title begins first play title and outside. Play Title, begins first play title, and outside. Play Title. begins first play title. and outside. Play Title; begins first play title; and outside.

11.4 Nested endings

There are too many cases I think to test them all. I’m testing to three levels of nesting. This is a Book Title including Book Title including Book Title and ending first one and outside. This is a Book Title including Book Title including Book Title, and ending first one, and outside. This is a Book Title including Book Title including Book Title. and ending first one. and outside. This is a Book Title including Book Title including Book Title; and ending first one; and outside. This is a Book Title including Book Title including Play Title and ending first one and outside. This is a Book Title including Book Title including Play Title, and ending first one, and outside. This is a Book Title including Book Title including Play Title. and ending first one. and outside. This is a Book Title including Book Title including Play Title; and ending first one; and outside. This is a Book Title including Play Title including Book Title and ending first one and outside. This is a Book Title including Play Title including Book Title, and ending first one, and outside. This is a Book Title including Play Title including Book Title. and ending first one. and outside. This is a Book Title including Play Title including Book Title; and ending first one; and outside.

This is a Book Title including Play Title including Play Title and ending first one and outside. This is a Book Title including Play Title including Play Title, and ending first one, and outside. This is a Book Title including Play Title including Play Title. and ending first one. and outside. This is a Book Title including Play Title including Play Title; and ending first one; and outside. This is a Play Title including Play Title including Play Title and ending first one and outside. This is a Play Title including Play Title including Play Title, and ending first one, and outside. This is a Play Title including Play Title including Play Title. and ending first one. and outside. This is a Play Title including Play Title including Play Title; and ending first one; and outside. This is a Play Title including Play Title including Book Title and ending first one and outside. This is a Play Title including Play Title including Book Title, and ending first one, and outside. This is a Play Title including Play Title including Book Title. and ending first one. and outside. This is a Play Title including Play Title including Book Title; and ending first one; and outside. This is a Play Title including Book Title including Play Title and ending first one and outside. This is a Play Title including Book Title including Play Title, and ending first one, and outside. This is a Play Title including Book Title including Play Title. and ending first one. and outside. This is a Play Title including Book Title including Play Title; and ending first one; and outside. This is a Play Title including Book Title including Book Title and ending first one and outside. This is a Play Title including Book Title including Book Title, and ending first one, and outside. This is a Play Title including Book Title including Book Title. and ending first one. and outside. This is a Play Title including Book Title including Book Title; and ending first one; and outside.

11.5 double and single nosuck

OS=open-single OD=open-double CS=close-single CD=close-double The following pairs of lines in medium weight roman should look identical. The line in typewriter font is the source text.

The following line in medium weight roman is what that source produces. The third line is what the second line ought to produce:

The word \WrapquotesDN{quoted} is quoted.

The word “quoted” is quoted. The word “quoted” is quoted.

The word \WrapquotesSN{scare} is in scare quotes.

The word ‘scare’ is in scare quotes. The word ‘scare’ is in scare quotes.

Nesting with no abuttment:

\WrapquotesDN{The \WrapquotesSN{quick} brown fox \WrapquotesDN{jumped} over the lazy dogs.}

“The ‘quick’ brown fox “jumped” over the lazy dogs.” “The ‘quick’ brown fox “jumped” over the lazy dogs.”

\WrapquotesSN{The \WrapquotesSN{quick} brown fox \WrapquotesDN{jumped} over the lazy dogs.}

‘The ‘quick’ brown fox “jumped” over the lazy dogs.’ ‘The ‘quick’ brown fox “jumped” over the laxy dogs.’

OS+OS, CD+CS:

\WrapquotesSN{\WrapquotesSN{The quick} brown fox jumped over the \WrapquotesDN{lazy dogs.}}

‘ ‘The quick’ brown fox jumped over the “lazy dogs.” ’ ‘ ‘The quick brown fox jumped over the “lazy dogs.” ’

OS+OD, CD+CS:

\WrapquotesSN{\WrapquotesDN{The quick} brown fox jumped over the \WrapquotesSN{lazy dogs.}}

‘ “The quick” brown fox jumped over the ‘lazy dogs.’ ’ ‘ “The quick” brown fox jumped over the ‘lazy dogs.’ ’

OD+OD, CS+CD:

\WrapquotesDN{\WrapquotesDN{The quick} brown fox jumped over the \WrapquotesSN{lazy dogs.}}

“ “The quick” brown fox jumped over the ‘lazy dogs.’ ” “ “The quick” brown fox jumped over the ‘lazy dogs.’ ”

OS+OD, CS+CS:

\WrapquotesSN{\WrapquotesDN{The quick} brown fox jumped over the \WrapquotesSN{lazy dogs.}}

‘ “The quick” brown fox jumped over the ‘lazy dogs.’ ’ ‘ “The quick” brown fox jumped over the ‘lazy dogs.’ ’

References

University of Chicago Press. 1993. The Chicago Manual of Style. 14th ed. Chicago: University of Chicago Press.

Index

Numbers written in italic refer to the page where the corresponding entry is described; numbers underlined refer to the code line of the definition; numbers in roman refer to the code lines where the entry is used.

Symbols

\, 195, 198, 201, 204, 207, 210, 213, 394, 401, 409, 423

\/ . . . . . . . . . . . . . . 259 \: . . . . . . . . . . 388–390 \@ . . . . . . . . . . 175, 443

\@break@tfor . . . . . . 342
\@doublequotes@false

. 29, 116, 131, \underline{141}

\@doublequotes@true

. 35, 120, 135, \underline{141}

\@firstoftwo . . . . . . 45
\@ifundefined . . . . . 11
\@look@nosuck@false

. . . . . . . . \underline{231}, 263

\@look@nosuck@true .

. . . . . . . . \underline{231}, 269 \@nil …. 42, 43, 53,

70, 73, 80, 97, 100

\@secondoftwo . . . . . 47

\@sptoken …. 259, 312

\@tempswafalse . 41, 331
\@tempswatrue . . . . .

59, 75, 86, 102, 341 \@tfor . . . . . . . . . . . 332

\@ti@sw@false . . . . . \underline{40}
\@ti@sw@true . . . . . . \underline{40}

\@whilesw . . . . . . 64, 91 \@xobeysp . . . . . . . . 259

\␣ . . . . . . . . . . . . . . 259

A

\aftergroup . . . 218,

221, 245, 248, 307, 311, 317, 383

\article . . . . . . . 4, \underline{447}

B

\begingroup . . . 113,

128, 162, 168, 174 \bgroup . . . . . . . . . . 259 \book . . . . . . . . . . 4, \underline{447}

C

\chaptertitle . . . . . 11

                          \check@icl . . . . . . . 366
                          \check@icr . . . . . . .

. 303, 366, 378, 381 \craft . . . . . . . . . 5, \underline{447} \csname . . . . . . . . . . 15

D

                          \DeclareOption . . 26, 32
                          \DeclareRobustCommand

. . . . . . . . . . . 360

                          \DeclareTextFontCommand

. . . . . . . . . . . 359 \def . 1–6, 27, 33, 54,

55, 81, 82, 117, 121, 132, 136, 311, 317, 359, 389

                          \defcommand . . . . . . . 305

\defn . . . . . . . . . . 4, \underline{433} \DefName . . . . . . . . . 378 \DGobbleM . . . . . . . . 266 \do . . . . . . . . . . . . . 337 \docdate . . . . . . . . \underline{1}, 5

                          \DoXPackageS . . . . . . . 2
                          \DoXUsepackagE . . . . . \underline{1}
                          \DTypeout . 108, 112,

115, 119, 130, 134, 143, 147, 152, 156, 161, 167, 173, 194, 197, 200, 203, 206, 209, 212, 234, 237, 240, 262, 265, 268, 271, 274, 281, 291, 300, 302, 304, 306, 310, 313, 316, 319, 322, 325, 328, 339, 344, 349, 352, 356, 372, 392, 393, 395, 399, 400, 402, 406, 408, 410, 413, 414, 420, 422, 424, 427, 428

E \E@car . . . . . . . . 73, 100 \E@cdr . . . . . . . . . 70, 97 \edef . . 7, 70, 73, 97, 100 \eExpand . . . . . . . . . 15 \egroup . . . . . . 259, 315 \else . 46, 57, 60, 67,

84, 87, 94, 118, 133, 196, 199, 202, 205, 208, 211, 214, 294, 309, 312, 315, 318, 321, 324, 327, 330, 343, 351, 363, 366, 378, 380, 412, 426

                                                   \endcsname . . . . . . . 15
                                                   \endgroup . 124, 139,

164, 170, 176, 309 \endinput . . . . . . . . 20

                                                   \essaytitle . . . . . . . 10
                                                   \ExecuteOptions . . . 38
                                                   \expandafter . . . 45,

47, 236, 270, 301, 332–335, 338, 340, 367, 389

F \fi 48, 64, 71, 76–78,

91, 98, 103–105, 122, 137, 215, 296, 346, 354, 355, 366, 368, 378, 382, 416, 430

\filedate . . . . . . . \underline{1}, 4 \fileinfo . . . . . . . \underline{1}, 2

                                                   \fileversion . . . . . \underline{1}, 3

\film . . . . . . . . . . . . . 9 \foreign . . . . . . . 3, \underline{433}

                                                   \foreignword . . . . 3, \underline{433}
                                                   \FrankenInfo . . . . . . 373
                                                   \futurelet . . . 189, 288

G

                                                   \g@addto@macro . . . .

. 394, 401, 409, 423 \gdef . . . . . . . . 181,

185, 224, 228, 251, 255, 282–284

H

\HaveECitationS . . . . \underline{1}

I

\if@doublequotes@ .

…. 114, 129, \underline{141}

\if@look@nosuck@ . .

…. \underline{231}, 407, 421

\if@tempswa . . . . 44, 348
\if@ti@sw@ . . . . . . . \underline{40}

\ifmmode . . . . . . . . . 361

\IfQuestionOrExclamation

. . . . . . . 7, \underline{40}, 109 \ifvmode . . 292, 366,

374, 376, 378, 379 \ifx 56, 58, 65, 74, 83,

85, 92, 101, 193, 196, 199, 202, 205, 208, 211, 292, 303, 309, 312, 315, 318, 321, 324, 327, 338

\In . . . . . . . . . . . . . 15

\initelyHavECitationS

. . . . . . . . . . . . . 3

\InputIfFileExists . . 1

J

\journal . . . . . . . 4, \underline{447}

\JusTLoaDInformatioN 13

L

\leavevmode . . . . . . . 364

\let 7, 28, 34, 61, 62, 68, 69, 88, 89, 95, 96, 293, 295, 314, 320, 323, 326, 329, 350, 353, 388, 390, 411, 415, 425, 429

\longpoem . . . . . . . . . 8

M

\makeatletter . . . . . 10
\makeatother . . . . . . 19
\manualref . . . . . . . 453

\meaning . . 237, 240, 265, 268, 271, 274, 291, 306, 339, 340, 344, 345, 349, 352, 356, 392, 393, 395, 399, 400, 402, 406, 410, 414, 420, 424, 428

\MessageBreak . 374–376

\music . . . . . . . . . 4, \underline{447}

N

\NeedsTeXFormat . . . 22
\newboolean 50, 141, 231
\newcommand . . . . 40,

107, 127, 142, 146, 151, 155, 160, 166, 172, 180, 184, 188, 191, 217, 220, 223, 227, 233, 244, 247, 250, 254, 261, 279, 287, 290, 299, 371, 391, 398, 405, 419, 435, 442

\newlet . . 8–11, 433,

434, 438, 445–455 \NewName . . . . . . . 53, 80

\nfss@text . . . . . . . 362
\nospacelist . . . . . .

…. 236, \underline{258}, 270

P

\phrase . . . . . . . . 3, \underline{433} \play . . . . . . . . . . 5, \underline{447}

\poemtitle . . . . . 5, \underline{447}

\PPOptArg . . . . . . . \underline{1}, 23

\ProcessOptions . . . 39
\ProvidesFile . . . . . . 6
\ProvidesPackage . . 23

R

\requirecommand . . . 258
\RequirePackage . . . 24
\ReserveCS . . . . . . .

25, 51, 52, 178, 387

\RestoreDoXVarS . . . 18

S

\SaveDoXVarS . . . . . . 14

\sc@t@a . . . . . 54, 56, 58, 61, 62, 65, 69, 70, 81, 83, 85, 88, 89, 92, 96, 97, 117, 121, 123, 132, 136, 138, 189, 193, 196, 199, 202, 205, 208, 211, 283, 349, 350, 388, 390, 411, 425 \sc@t@b . . 55, 58, 74,

82, 85, 101, 284, 352, 353, 415, 429

\sc@t@c 237, 240, 265, 266, 268, 271, 274, 288, 291, 292, 301, 303, 309, 312, 315, 318, 321, 324, 327, 338, 339, 344

\sc@t@d … 293, 295, 297, 305, 311, 314, 317, 320, 323, 326, 329, 350, 353, 356, 357, 375, 385, 411, 415, 425, 429

\sc@t@e . . . . . . 282, 336 \sc@t@f 181, 185, 192,

224, 228, 237, 238, 240, 241, 251, 255, 265, 266, 268, 271, 272, 274, 275, 393–395, 400– 402, 409, 410, 414, 423, 424, 428

\sc@t@g 333, 338, 340, 345

\sc@toks@a . . . . . . .

. 280, 281, 301, 302

\ShortEmpty 56, 65, 83, 92

\slash . . . . . . . . . . . 259 \space . . . . . . 8, 241,

259, 275, 374–376 \species . . . . . . . 5, \underline{447}

\storytitle . . . . . . . \underline{447}

\string … 374–376, 408, 413, 422, 427

T

\term . . . . . . . . . . 4, \underline{433}

\text@command . . . . . 365
\textitswitch . . . 8,

9, 433, 434, 445, 447–449, 453–455

\textquotedblleft . 181
\textquotedblright .

. 228, 251, 394, 409

\textquoteleft 185, 443
\textquoteright 224,

255, 401, 423, 443

\textslswitch . . . 7, 446

\the . . . . . . . . . 281, 302

\ti@@openquote 189, 191

\ti@checkfor@e . . . . \underline{40}
\ti@checkfor@q . . . . \underline{40}
\ti@close@double@@nosuck

. . . . . . . . . . . \underline{217}

\ti@close@double@@suck

. . . . . . . . 245, \underline{250}

\ti@close@double@nosuck

…. 132, 169, \underline{217}, 324, 408, 413

\ti@close@double@suck

…. 117, \underline{244}, 318

\ti@close@quote@nosuck

…. 225, 229, \underline{231}

\ti@close@quote@suck

…. 252, 256, \underline{261}

\ti@close@single@@nosuck

. . . . . . . . . . . \underline{217}

\ti@close@single@@suck

. . . . . . . . 248, \underline{250}

\ti@close@single@nosuck

. 136, 163, 175, \underline{217}, 327, 422, 427

\ti@close@single@suck

…. 121, \underline{244}, 321

\ti@domelater . . \underline{25}, 179

\ti@open@double . . .

. 117, 132, 169, \underline{180}

\ti@open@single 121,

136, 163, 175, \underline{180}

\ti@openquote . . . . . \underline{180}

\ti@prev . . . . . . . . . \underline{40}

\ti@prev@prev . . . . . \underline{40}
\ti@q@@ifnch . . . . . . \underline{279}
\ti@q@check . . . . . . .

. \underline{279}, 383, 389, 396, 403, 417, 431

\ti@q@handle@double@nosuck

. . . . . . . . 326, \underline{405}

\ti@q@handle@double@suck

. . . . . . . . 320, \underline{391}

\ti@q@handle@ifvmode

. . . . . . . . 293, \underline{359}

\ti@q@handle@single@nosuck

. . . . . . . . 329, \underline{405}

\ti@q@handle@single@suck

. . . . . . . . 323, \underline{391}

\ti@q@handle@space .

. . . . . . . . 314, \underline{387}

\ti@q@ifnch . . . . . . . \underline{279}
\ti@q@ifnextcharin .

. 236, 264, 270, \underline{279}

                          \ti@wrapquotes@nosuck

…. \underline{107}, 153, 158

                          \ti@wrapquotes@suck

…. \underline{107}, 144, 149

                          \ToggleBoolean 148, 157

U

                          \UndefineCS . . . . . . . 13
                          \usepackage . . . . . . . 16

W \word . . . . . . . . 3, 7, \underline{433}

                          \Wrapquotes 6, 10, 11,

28, 34, \underline{141}, 450–452

                          \WrapquotesDN 6, \underline{141}, 208
                          \WrapquotesIN 6, \underline{141}, 199
                          \WrapquotesIS 6, \underline{141}, 202
                          \WrapquotesNN . . . . .

… 6, 28, \underline{141}, 196

                          \WrapquotesNS . . . . .

… 6, 34, \underline{141}, 193

                          \WrapquotesSK . . . . .

. . 6, \underline{141}, 211, 438

                          \WrapquotesSN 6, \underline{141}, 205